OpenStudy (anonymous):

We wish to differentiate sin^2(x) . For that, we derivate D(sin2(x))=D(sin(x)sin(x)) = __________= 2sin(x)cos(x)=_____________ Therefore D(sin2(x)=sin(2x). I'm supposed to fill in the blanks

OpenStudy (slaaibak):

First blank space: D(sin(x)sin(x)) = D(sin(x)) * sin(x) + sin(x)*D(sin(x)) = cos(x)sin(x) + sin(x)cos(x) = 2sin(x)cos(x) = sin(2x)

OpenStudy (slaaibak):

Which part is giving you problems?

OpenStudy (anonymous):

nvm >,<

OpenStudy (anonymous):

how does cos(x)sin(x)+sin(x)cos(x) = 2 sin(x) cos(x)??

OpenStudy (anonymous):

2 ( sin(x)+cos(x))??

OpenStudy (slaaibak):

Nope. Do you know that xy = yx? Then further, do you know that xy + yx = xy + xy = xy(1 + 1) = 2xy ?

OpenStudy (anonymous):

its a property...

OpenStudy (slaaibak):

? cos(x)sin(x {xy})+sin(x)cos(x) {yx} apply the same logic

OpenStudy (anonymous):

lmao omg thanks